OpenStudy (anonymous):

Find the balance in the account: $1,600 principle earning 7%, compounded semi-annually, after 33 years.

OpenStudy (ybarrap):

balance = 1,600 * (1.07)^(33) That's it! http://www.wolframalpha.com/input/?i=+1%2C600+*+%281.07%29%5E%2833%29

OpenStudy (tkhunny):

Or, with the semi-annual compounding: \(1600\cdot\left(1+\dfrac{0.07}{2}\right)^{33*2}\)

OpenStudy (anonymous):

That's so confusing.

OpenStudy (tkhunny):

Annual Compunding (1 per year) \(1600\cdot\left(1 + \dfrac{0.07}{1}\right)^{33\cdot 1}\) Semi-Annual Compunding (2 per year) \(1600\cdot\left(1 + \dfrac{0.07}{2}\right)^{33\cdot 2}\) Quarterly Compunding (4 per year) \(1600\cdot\left(1 + \dfrac{0.07}{4}\right)^{33\cdot 4}\) Monthly Compunding (12 per year) \(1600\cdot\left(1 + \dfrac{0.07}{12}\right)^{33\cdot 12}\) Weekly Compunding (52 per year) \(1600\cdot\left(1 + \dfrac{0.07}{52}\right)^{33\cdot 52}\) Daily Compunding (365 per year) \(1600\cdot\left(1 + \dfrac{0.07}{365}\right)^{33\cdot 365}\) It all relates quite nicely.

OpenStudy (anonymous):

= 1600*(1.035)^66

OpenStudy (ybarrap):

@tkhunny is correct. I missed the semi-annual part of the problem